integrate sinx/sin4x

Grade:

1 Answers

KADAMBALA VAMSI
10 years ago

(sinx)/(sin4x) dx

=∫1/4cosxcos2x  dx

=1/4 ∫cosx dx /cos2xcos2x

==1/4 ∫cosx dx /(1-sin2x)(1-2sin2x)

let sinx =t

   cosx dx =dt

=1/4 ∫dt /(1-t2)(1-2t2)

use partial fraction

=-1/4 ∫dt /(1-t2)   + 1/8 ∫dt /(1-2t2)

now onwards it is easily calculated